A force of 3.0 N acts through a distance of 12m in the direction of the force. Find the work done.


1
Expert's answer
2022-05-18T15:28:01-0400
"W=Fd\cos\theta\\
=\rm 3.0\: N\times 12\: m\times \cos0^\circ=36\: N"